Given the following data, find the lower bound of confidence interval .Your answer must be in two decimal places.

Sample size 100, Mean = 80, Standard Deviation = 20, Confidence coefficient 1.645 -
Answer #1

Lower bound on confidence interval :=

= mean - confidence coefficient* standard deviation / sqrt(sample size)

= 80 - 1.645*20/sqrt (100) = 80 - 3.29 = 76.71